# Thermodynamic Entropic Uncertainty Relation

### v1

@@ Abstract
Thermodynamic uncertainty relations reveal a fundamental trade-off between the precision of trajectory observables and entropy production, where the uncertainty in observables is quantified by their variance. In the context of information theory, uncertainty is often evaluated in terms of Shannon entropy, but it remains unclear whether there is a quantitative relation between Shannon entropy of the observables and entropy production in stochastic thermodynamics. In this Letter, we show that uncertainty relations can be formulated with observable Shannon entropy and entropy production. We introduce symmetry entropy, an entropy measure that quantifies the symmetry of the observable distribution, and demonstrate that a greater asymmetry in the observable distribution demands higher entropy production. This Letter elucidates the role of Shannon entropy of observables within stochastic thermodynamics, thereby establishing a foundation for deriving new uncertainty relations.
Introduction.—Quantum mechanics operates in ways that are fundamentally different from classical physics. The Heisenberg uncertainty relation, proposed by Heisenberg in 1927 [1], captures the unique nature of quantum mechanics through a single inequality, rendering the inability to determine position and momentum precisely. Robertson [2] generalized the Heisenberg uncertainty relation so that the relation can incorporate observables other than position or momentum. Given a quantum state |$\ket{\psi}$ and observables $A$ and $B$, the Robertson uncertainty relation states
$\mathrm{Var}[A]\mathrm{Var}[B]\geq\frac{1}{4}|\braket{\psi}{[A,B]}{\psi}|^{2},$ (1)
where [,][$\bullet$,$\bullet$][ ∙ , ∙ ] is the commutator and $\mathrm{Var}[A]$ is the variance of $A$ with respect to |$\ket{\psi}$. The Robertson uncertainty relation is recognized for its fundamental role in deriving various other relations, including the quantum speed limit [3, 4, 5]. Although the Robertson uncertainty relation given by Eq. (1) quantifies the uncertainty of the observables via their variance, uncertainty is often evaluated using the Shannon entropy in the context of information theory. It is therefore natural to expect uncertainty relations that incorporate the Shannon entropy. Indeed, Ref. [6] showed that the uncertainty relation involving the entropy of position and momentum holds in quantum mechanics. Several generalizations and extensions have been proposed for the entropic uncertainty relation [7] and the most well known instance is Maassen-Uffink relation [8, 9]:
$H[A]+H[B]\geq\ln\frac{1}{c}.$ (2)
Here, $H[A]$ and $H[B]$ denotes the Shannon entropy of measurement outputs of $A$ and $B$, respectively, and $c$ denotes the maximum overlap between two eigenvectors of $A$ and $B$. The entropic uncertainty relation is crucial not only for understanding the nature of quantum mechanics but also plays a vital role in quantum cryptography, especially in quantum key distribution protocols.
In recent years, it has become clear that uncertainty relations are prevalent in stochastic thermodynamics [10, 11]. In particular, the thermodynamic uncertainty relation [12, 13] indicates a trade-off between entropy production and the relative variance of the current observable. For a stochastic thermodynamic system in the steady state and a trajectory observable $F$, the following relation holds:
@@
where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ denote the expectation and the variance of $F$, respectively, and $\Sigma$ is the entropy production within the time interval of interest. Equation (3) suggests that achieving greater precision requires increased entropy production and signifies a no-free lunch in thermodynamic systems. The thermodynamic uncertainty relation in Eq. (3) resembles the Robertson uncertainty relation in that the uncertainty is evaluated with the variance. In fact, certain types of thermodynamic uncertainty relations are known to actually be derived from the Robertson uncertainty relation [14, 15]. Given this context, a simple question arises: Does an entropic uncertainty relation hold in stochastic thermodynamics? In this paper, we confirm that this conjecture is correct and demonstrate that there is an uncertainty relation between entropy production and the symmetry entropy [cf. Eq. (12)], which quantifies the extent of symmetry of observable distributions. Specifically, we show that the sum of the entropy production and the symmetry entropy should be no less than $\ln 2$ [cf. Eqs. (14) and (16)]. In other words, the asymmetry of the observable probability distribution, as quantified by entropy, requires that the entropy production be at least equal to this measure of asymmetry. We validate the uncertainty relation with a biased random walk, which is often used to model biomolecular processes. The trade-off relationship in thermodynamic cost has been extensively studied, particularly in terms of thermodynamic uncertainty relations [12, 13, 16, 17, 18, 19, 20, 21, 22, 23] using the variance of observable quantities and the speed limits using the distance between states [24, 25, 26, 22, 27, 28, 29]. This study demonstrates a trade-off between the entropy of observables and the entropy production, which is expected to lead to the derivation of other trade-off relations.
@@
Figure 1: Examples of the symmetry entropy $\Lambda[P(\Phi)]$. Horizontal axes denote values of $\Phi$ (left column) and $|\Phi|$ (right column). The values of $\Phi$ always form pairs; $-1$ and $1$, and $-2$ and $2$ are two pairs in the examples. Vertical axes denote probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). (a) Symmetric distribution where probabilities of all of the pairs are identical. $P(|\Phi|)$ is different from the original distribution $P(\Phi)$ and thus $\Lambda[P(\Phi)]$ is $\ln 2$. (b) Asymmetric distribution where probabilities of all of the pairs are fully biased. $P(|\Phi|)$ and $P(\Phi)$ are identical and thus $\Lambda[P(\Phi)]$ is 00. (c) Asymmetric distribution where probabilities of all of the pairs are fully biased. The distributions $P(|\Phi|)$ and $P(\Phi)$ are different, but they effectively become the same if the labels for the pair $-1$ and $1$ are swapped. Therefore, they are essentially the same distribution. This results in $\Lambda[P(\Phi)]=0$.
@@
It is known that the Kullback-Leibler divergence is non-negative. Moreover, the Kullback-Leibler divergence satisfies monotonicity. Consider a transformation that maps the original random variables $X$ and $Y$ to new random variables $\tilde{X}$ and $\tilde{Y}$, respectively. Then the following monotonicity relation holds:
@@
where $P(\tilde{X})$ and $P(\tilde{Y})$ are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$. The Kullback-Leibler divergence is a measure that quantifies the distance between two probability distributions. The monotonicity shown by Eq. (7) implies that no matter what transformation is applied to random variables, the distance between the two probability distributions cannot be increased.
Having introduced basic concepts of the divergence, we move to consideration of stochastic thermodynamic systems. Stochastic thermodynamics considers processes whose state change is described by a stochastic process. Let $\Gamma$ be a stochastic trajectory of the process and $\Gamma^{\dagger}$ be its time reversal. Moreover, we can define the probability of measuring $\Gamma$, which is denoted by $\mathcal{P}(\Gamma)$. Assuming the local detailed balance, it is known that the entropy production under the steady-state condition is defined by the Kullback-Leibler divergence:
@@
Note that the expression of Eq. (8) also holds for continuous processes such as Langevin dynamics. Consider an observable $\Phi(\Gamma)$, which is a function of the trajectory $\Gamma$. Here, we assume that $\Phi(\Gamma)$ is anti-symmetric under the time reversal:
@@
For example, $\Phi(\Gamma)$ represents a thermodynamic current. Important thermodynamic quantities, such as stochastic dissipated heat or displacement, are expressed by $\Phi(\Gamma)$. Here, we initially assume that $\Phi(\Gamma)$ takes values in a countable set, that is, the probability distribution $P(\Phi)$ is discrete. However, all the results below hold for the continuous case as well.
Figure 2: Illustration of the example. (a) Biased random walk with right and left transition probabilities are $w_{+}$ and $w_{-}$, respectively. (b) Illustration of F1-ATPase. This biased random walk can model molecular motors such as F1-ATPase, where the motor uses ATP hydrolysis to rotate discretely in steps of $32\pi/$ 3 in the counter-clockwise direction.
Results.—We derive the thermodynamic entropic uncertainty relation, which is the main result of this Letter. Using the monotonicity of Kullback-Leibler divergence [Eq. (7)] and the time reversal property of $\Phi$ [Eq. (9)], we have
$\Sigma$ $\geq D[P(\Phi)\|P(-\Phi)]$
$=-H[P(\Phi)]+C[P(\Phi),P(-\Phi)].$ (10)
Here, the cross entropy is evaluated as $C[P(\Phi),P(-\Phi)]=-\sum_{\phi}P(\Phi=\phi)\ln P(\Phi=-\phi)$. Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$. An important observable is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term. To simplify, let us consider that the observable $\Phi(\Gamma)$ does not include 00, whose condition is met for several problem settings. For instance, we may consider binary classification using trajectories of stochasitc processes. In this case, the observable $\Phi(\Gamma)$ does not include 00. $\Phi(\Gamma)=0$ should be handled separately because $\Phi(\Gamma)=\Phi(\Gamma^{\dagger})=0$ from Eq. (9), showing that the observable is invariant under the time reversal. Later, we will consider the case where $\Phi(\Gamma)=0$ is included. Specifically, it can be proved that the following relation holds:
$C[P(\Phi),P(-\Phi)]\geq\ln 2+H[P(|\Phi|)],$ (11)
whose derivation is shown in Appendix A. In Eq. (11), $P(|\Phi|)$ is the probability distribution of $|\Phi|$, which is the absolute random variable of $\Phi$. Let $\Lambda[P(\Phi)]$ be
@@
which is the entropy difference between the original distribution $P(\Phi)$ and its absolute valued distribution $P(|\Phi|)$. Here, we call $\Lambda[P(\Phi)]$ as symmetry entropy. The symmetry entropy $\Lambda[P(\Phi)]$ quantifies the extent of symmetry of $P(\Phi)$. Figure 1 depicts examples of values of the symmetry entropy, where the horizontal axes are values of $\Phi$ (left column) and $|\Phi|$ (right column) and the vertical axes are $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). From the condition of time reversal [Eq. (9)], if $\Phi=a$ exists ($a>0$), then $\Phi=-a$ also exists, which is regarded as a pair of the observable. Suppose that the distribution $P(\Phi)$ is symmetric for all of the pairs, as depicted in Fig. 1(a). For this case, $P(|\Phi|)$ is very different from the original distribution $P(\Phi)$, which results in $\Lambda[P(\Phi)]$ being $\ln 2$. In contrast, for an asymmetric distribution illustrated in Fig. 1(b), $P(\Phi)$ and $P(|\Phi|)$ are the same, and thus $\Lambda[P(\Phi)]$ reduces to 00. To be more specific, the symmetry entropy quantifies how biased the probability distributions of pairs of observables are. When there is a strong bias in the probabilities of the pairs, the value of $\Lambda[P(\Phi)]$ becomes small. For example, $\Lambda[P(\Phi)]$ becomes 00 not only in cases where $P(\Phi)$ and $P(|\Phi|)$ are the same, as shown in Fig. 1(b), but also in cases like Fig. 1(c). It can be shown that
@@
whose proof is provided in Appendix B. In Eq. (13), $\Lambda[P(\Phi)]$ being 00 and $\ln 2$ corresponds to asymmetric and symmetric distributions, respectively. Substituting Eq. (11) into Eq. (10), we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
@@
Equation (14) is the main result of this study and referred to as thermodynamic entropic uncertainty relation. The right-hand side of Eq. (14) quantifies the asymmetry of probability distribution $P(\Phi)$. Therefore, Eq. (14) shows that, for arbitrary observable $\Phi$ satisfying the time reversal condition [Eq. (9)], the system requires the entropy production no less than $\ln 2-\Lambda[P(\Phi)]$. The trade-off between entropy production and observable asymmetry parallels traditional thermodynamic uncertainty relations, which illustrate a trade-off between the variance of observables and entropy production. The thermodynamic uncertainty relations consider the relative variance $\mathrm{Var}[\Phi]/\mathbb{E}[\Phi]^{2}$. In a sense, using the relative variance can also be seen as quantifying the asymmetry of $P(\Phi)$; when the variance is smaller and the expectation is greater, the probability distribution $P(\Phi)$ is more asymmetric with respect to $\Phi=0$. Since $\ln 2-\Lambda[P(\Phi)]\geq 0$ in Eq. (14) due to Eq. (13), Eq. (14) can be regarded as a refinement of the second law using the entropy of the observable. There are some advantages in employing the entropy instead of the variance. When dealing with the variance, the observable must yield a real number. However, in cases where the observable consists of classifications such as “success” and “failure”, the variance is not suitable. Even in such cases, the Shannon entropy is well defined, indicating that Eq. (14) can be applied.
Thus far, we assumed that $\Phi(\Gamma)$ takes values in a countable set, that is, $P(\Phi)$ is a discrete distribution. Consider the case where $\Phi(\Gamma)$ produces continuous values, where the summation should be replaced by the integration. When considering a continuous distribution, the notable difference is that the differential entropy may take negative values. However, the existence of negative values is not problematic, as such negative values are offset in $H[P(\Phi)]-H[P(|\Phi|)]$. Therefore, the main result of Eq. (14) is well defined for the continuous case as well.
The derivation of Eq. (14) assumed that the observable $\Phi(\Gamma)$ does not include $\Phi(\Gamma)=0$. It is straightforward to extend the result to the case where the observable includes $\Phi(\Gamma)=0$. Specifically, when $\Phi(\Gamma)=0$ is included, the range of $\Lambda[P(\Phi)]$ is modified as follows:
@@
which includes Eq. (14) as the specific case $P(\Phi=0)=0$. As long as the entropy production is given by Eq. (8), Eq. (16) holds for arbitrary observable $\Phi(\Gamma)$ satisfying Eq. (9). In Eq. (16), $\Phi=0$ plays a special role. When considering the process of doing nothing, the observable $\Phi$ is always 00 implying $P(\Phi=0)=1$. For such empty dynamics, the entropy production is 00, showing that both sides of Eq. (16) equal 00 and thus the inequality becomes equality. This consideration shows that $\Phi=0$ should be handled separately. However, note that this exceptional handling of $\Phi=0$ is not limited to Eq. (16). In the conventional thermodynamic uncertainty relation [Eq. (3)], when the expectation of the current vanishes, the inequality is ill-defined. When $P(\Phi)$ is a probability density that is smooth around 00, the measure becomes $P(\Phi=0)=0$. Only when $P(\Phi)$ includes the delta-peaked contribution at $\Phi=0$, we use Eq. (16)
@@
Let us comment on the relation between Eq. (16) and the Landauer principle [30]. When considering a process that resets to one state from a state that exists with equal probability in two states, the following relation is obtained.
$\Delta S_{m}\geq\ln 2,$ (17)

### v3

@@ Abstract
Thermodynamic uncertainty relations reveal a fundamental trade-off between the precision of a trajectory observable and entropy production, where the uncertainty of the observable is quantified by its variance. In information theory, Shannon entropy is a common measure of uncertainty. However, a clear quantitative relationship between the Shannon entropy of an observable and the entropy production in stochastic thermodynamics remains to be established. In this Letter, we show that an uncertainty relation can be formulated in terms of the Shannon entropy of an observable and the entropy production. We introduce symmetry entropy, an entropy measure that quantifies the symmetry of the observable distribution, and demonstrate that a greater asymmetry in the observable distribution requires higher entropy production. Specifically, we establish that the sum of the entropy production and the symmetry entropy cannot be less than $\ln 2$. As a corollary, we also prove that the sum of the entropy production and the Shannon entropy of the observable is no less than $\ln 2$. As an application, we demonstrate our relation in the diffusion decision model, revealing a fundamental trade-off between decision accuracy and entropy production in stochastic decision-making processes.
Introduction.— The thermodynamic uncertainty relation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] expresses a universal trade-off between the entropy production and the relative variance of time-integrated observables for systems operating far from equilibrium. For a steady-state stochastic thermodynamic system and a trajectory observable $F$, the thermodynamic uncertainty relation states [1, 2]
@@
where $\mathbb{E}[F]$ and $\mathrm{Var}[F]$ are the expectation value and variance of $F$, respectively, and $\Sigma$ denotes the entropy production during the observation interval. The thermodynamic uncertainty relation implies that achieving higher precision inevitably requires greater entropy production–a manifestation of the “no free lunch” principle in thermodynamics. The thermodynamic uncertainty relation given by Eq. (1) quantifies the uncertainty of the observables via their variance; however, uncertainty is often evaluated using the Shannon entropy in the context of information theory. Consider a thermodynamic decision-making process that yields a categorical output, such as a “yes” or “no” label. Since the output is categorical rather than numerical, we cannot use variance to quantify its uncertainty. However, even in such cases, we can still calculate the Shannon entropy. Therefore, to analyze the tradeoff between the accuracy of the categorical outputs and the entropy production, a thermodynamic uncertainty relation based on Shannon entropy is required.
The issue of how to quantify uncertainty has been actively studied in quantum mechanics. In the Heisenberg-Robertson uncertainty relation [11, 12], the uncertainty in the outcomes of two observables is quantified through their variances. Later, Hirschman showed that the variance can be replaced by the Shannon entropy [13, 14]. Maassen and Uffink [15] derived a precise statement of this entropic uncertainty relation [16]: the sum of the entropies of the two observables cannot fall below a bound determined by the extent of their eigenstate overlap. The entropic uncertainty relation plays a fundamental role in quantum cryptography, especially in quantum key distribution protocols [17].
@@
In this Letter, in the context of stochastic thermodynamics, we show that there is an uncertainty relation between entropy production and the symmetry entropy [cf. Eq. (8)], which quantifies the extent of symmetry of the observable’s distribution. Specifically, we show that the sum of the entropy production and the symmetry entropy of the observable should be no less than $\ln 2$ [cf. Eqs. (10) and (14)]. In other words, the asymmetry of the observable probability distribution, as quantified by entropy, requires that the entropy production be at least equal to this measure of asymmetry. As a corollary of the result, we also show that the sum of the entropy production and the Shannon entropy of the observable should be no less than $\ln 2$ [Eq. (12)]. We apply the derived relation to the diffusion decision model, illustrating a trade-off between decision accuracy and entropy production in stochastic decision-making processes. The trade-off relation in thermodynamic cost has been actively studied, particularly in terms of thermodynamic uncertainty relations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] using the variance of observable quantities and speed limits using the distance between states [18, 19, 20, 9, 21, 22, 23]. This study demonstrates a trade-off between the entropy of the observable and the entropy production, which is expected to lead to the derivation of other trade-off relations.
Figure 1: Conceptual representation of the thermodynamic entropic uncertainty relation. (a) Stochastic thermodynamic process. The thermodynamic entropic uncertainty relation considers a stochastic process, where each state transition is random. (b) Trajectory of the stochastic process shown in (a). $\Gamma$ denotes time evolution of a realization of the process. $\Gamma^{\dagger}$ is the time-reversal of $\Gamma$. (c) Probability distribution of observable $\Phi(\Gamma)$. $\Phi(\Gamma)$ is arbitrary, provided it satisfies the time-reversal antisymmetric property [Eq. (7)].
@@
Figure 2: Examples of the symmetry entropy $\Lambda[P(\Phi)]$. Horizontal axes denote values of $\Phi$ (left column) and $|\Phi|$ (right column). The values of $\Phi$ always form pairs; $-1$ and $1$, and $-2$ and $2$ are two pairs in the examples. Vertical axes denote probability distribution $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). (a) Distribution symmetric with respect to $\Phi=0$ where probabilities of all of the pairs are identical. $P(|\Phi|)$ is different from the original distribution $P(\Phi)$ and thus $\Lambda[P(\Phi)]$ is $\ln 2$. (b) Asymmetric distribution where probabilities of all of the pairs are fully biased. $P(|\Phi|)$ and $P(\Phi)$ are identical and thus $\Lambda[P(\Phi)]$ is 00. (c) Asymmetric distribution where probabilities of all of the pairs are fully biased. The distributions $P(|\Phi|)$ and $P(\Phi)$ are different, but they effectively become the same if the labels for the pair $-1$ and $1$ are swapped. Therefore, they are essentially the same distribution. This results in $\Lambda[P(\Phi)]=0$.
@@
It is well known that the Kullback-Leibler divergence is non-negative. Moreover, the Kullback-Leibler divergence satisfies monotonicity. Consider a transformation that maps the original random variables $X$ and $Y$ to new random variables $\tilde{X}$ and $\tilde{Y}$, respectively. Then the following monotonicity relation holds:
@@
where $P(\tilde{X})$ and $P(\tilde{Y})$ are probability distributions of the transformed variables $\tilde{X}$ and $\tilde{Y}$, respectively.
Having introduced basic concepts of the divergence, we move to consider stochastic thermodynamic systems. Stochastic thermodynamics considers processes whose state change is described by a stochastic process (Fig. 1(a)). Let $\Gamma$ be a stochastic trajectory of the process and $\Gamma^{\dagger}$ be its time reversal (Fig. 1(b)). Moreover, we can define the probability of measuring $\Gamma$, which is denoted by $\mathcal{P}(\Gamma)$. Assuming the local detailed balance, it is known that the entropy production under the steady-state condition is defined by the Kullback-Leibler divergence [24, 25]:
@@
Note that the expression of Eq. (6) also holds for continuous processes such as Langevin dynamics. Consider an observable $\Phi(\Gamma)$, which is a function of the trajectory $\Gamma$. Here, we assume that $\Phi(\Gamma)$ is antisymmetric under the time reversal:
@@
For example, $\Phi(\Gamma)$ represents a thermodynamic current. Important thermodynamic quantities, such as stochastic dissipated heat or displacement, are expressed by $\Phi(\Gamma)$. In the supplementary material [26], we provide a generalized formulation where the time-reversal antisymmetric condition [Eq. (7)] is relaxed. Initially, we assume that $\Phi(\Gamma)$ takes values in a countable set, that is, the probability distribution $P(\Phi)$ is discrete (Fig. 1(c)). However, most of the results below hold for the continuous case as well.
Figure 3: (a) Description of the diffusion decision model. The green trajectory illustrates a realization that exceeds the threshold $\theta$, resulting in a decision of $+1$ (“yes” class). Conversely, the purple trajectory represents a realization that falls below $-\theta$, leading to a decision of $-1$ (“no” class). (b) and (c) Entropies as a function of parameter $f$. In both figures, the blue dashed line shows the entropy production $\Sigma$, and the purple dot-dashed line represents $H[P(\Phi)]$, the entropy of the distribution $P(\Phi)$. Their sum, $\Sigma+H[P(\Phi)]$, is also depicted by the red solid line. The orange dotted line represents $\ln 2$ and $(1-P(0))\ln 2$ in (a) and (b), respectively, which serves as the lower bound for $\Sigma+H[P(\Phi)]$ according to Eq. (15). These plots are shown for $\theta=0$ in (b) and $\theta=0.5$ in (c), with other parameters being $A=1$ and $\tau=10$.
Results.—We derive the thermodynamic entropic uncertainty relation, which is the main result of this Letter. We first introduce an entropic measure which quantifies the observable of a trajectory $\Gamma$. Let $\Lambda[P(\Phi)]$ be
@@
which is the entropy difference between the original distribution $P(\Phi)$ and its absolute valued distribution $P(|\Phi|)$. Here, we call $\Lambda[P(\Phi)]$ symmetry entropy. The symmetry entropy $\Lambda[P(\Phi)]$ quantifies the extent of symmetry of $P(\Phi)$. Intuitively speaking, the difference $H[P(\Phi)]-H[P(|\Phi|)]$ quantifies how much “sign” information remains in the distribution $P(\Phi)$ once its magnitude $P(|\Phi|)$ is known. If $P(\Phi)$ is perfectly symmetric with respect to $\Phi=0$, knowing $P(|\Phi|)$ tells us nothing about the sign, so $\Lambda[P(\Phi)]$ reaches its maximum of $\ln 2$ (one bit). In contrast, if $P(\Phi)$ is fully biased in one direction, the sign is already determined by $P(|\Phi|)$ and $\Lambda[P(\Phi)]$ vanishes. Figure 2 illustrates examples of the symmetry entropy, with the horizontal axes representing the values of $\Phi$ (left column) and $|\Phi|$ (right column), and the vertical axes corresponding to $P(\Phi)$ (left column) and $P(|\Phi|)$ (right column). From the time-reversal antisymmetric condition [Eq. (7)], if $\Phi=a$ exists ($a>0$), then $\Phi=-a$ also exists, which is regarded as a pair of the observable. Suppose that the distribution $P(\Phi)$ is symmetric with respect to $\Phi=0$ for all of the pairs, as depicted in Fig. 2(a). For this case, $P(|\Phi|)$ is very different from the original distribution $P(\Phi)$, which results in $\Lambda[P(\Phi)]=\ln 2$. In contrast, for an asymmetric distribution illustrated in Fig. 2(b), $P(\Phi)$ and $P(|\Phi|)$ are the same, and thus $\Lambda[P(\Phi)]$ reduces to 00. Similarly, $\Lambda[P(\Phi)]$ is 00 for the bias pattern represented in Fig. 2(c). When $\Phi(\Gamma)$ is a binary function (outputs two categories), we have $\Lambda[P(\Phi)]=H[P(\Phi)]$.
For simplicity, let us consider the case where the observable $\Phi(\Gamma)$ does not take the value 00, a condition which holds in several problem settings. For instance, we may consider binary classification based on trajectories of stochastic processes. In this case, the observable $\Phi(\Gamma)$ does not include 00. $\Phi(\Gamma)=0$ should be handled separately because $\Phi(\Gamma)=\Phi(\Gamma^{\dagger})=0$ from Eq. (7), showing that the observable is invariant under the time reversal. Later, we will consider the case where $\Phi(\Gamma)=0$ is included. It can be shown that
@@
whose proof is provided in the End Matter. In Eq. (9), $\Lambda[P(\Phi)]$ being 00 and $\ln 2$ corresponds to asymmetric and symmetric distributions, respectively. Using the symmetry entropy $\Lambda[P(\Phi)]$, we obtain the trade-off between the entropy production and the asymmetry of $P(\Phi)$ quantified by entropy:
@@
Equation (10) is the main result of this study and referred to as thermodynamic entropic uncertainty relation. The derivation is shown in the End Matter. The right-hand side of Eq. (10) quantifies the asymmetry of the probability distribution $P(\Phi)$. Equation (10) demonstrates that for any observable $\Phi$ that satisfies the time-reversal antisymmetric condition specified in Eq. (7), the entropy production must be at least $\ln 2-\Lambda[P(\Phi)]$. The trade-off between entropy production and observable asymmetry parallels traditional thermodynamic uncertainty relations, which illustrate a trade-off between the relative variance of observables and entropy production. The thermodynamic uncertainty relations consider the relative variance $\mathrm{Var}[\Phi]/\mathbb{E}[\Phi]^{2}$. In a sense, using the relative variance can also be seen as quantifying the asymmetry of $P(\Phi)$; when the variance is smaller and the expectation is greater, the probability distribution $P(\Phi)$ is more asymmetric with respect to $\Phi=0$. Since $\ln 2-\Lambda[P(\Phi)]\geq 0$ in Eq. (10) due to Eq. (9), Eq. (10) can be regarded as a refinement of the second law, employing the entropy of the observable. There are some advantages in employing the entropy instead of the variance. When the observable is a categorical variable that outputs labels such as “yes” or “no”, which is the case in classification tasks, the use of variance is not appropriate. Even in such cases, the Shannon entropy is well defined, indicating that Eq. (10) can be applied. The right-hand side of Eq. (10) can also be interpreted as a divergence. If we denote the Jensen-Shannon divergence by $\mathrm{JS}[P(X)\|P(Y)]$, then
$\mathrm{JS}[P(\Phi)\|P(-\Phi)]=\ln 2-\Lambda[P(\Phi)],$ (11)
the derivation of which is shown in the End Matter. Equation (10) provides a trade-off between the symmetry entropy and the entropy production. For a discrete probability distribution, $H[P(|\Phi|)]\geq 0$ holds. Therefore, $H[P(\Phi)]\geq\Lambda[P(\Phi)]$ and thus the following bound also holds:
$\Sigma\geq\ln 2-H[P(\Phi)],$ (12)
which purely relates the entropy production $\Sigma$ and the Shannon entropy of the observable $\Phi$. The right-hand side of Eq. (12) may not always be non-negative. Although the form of Eq. (12) is more appealing in terms of physical interpretation, the bound is weaker. When the observable $\Phi(\Gamma)$ is a binary function, the right-hand side of Eq. (12) becomes non-negative as $H[P(\Phi)]=\Lambda[P(\Phi)]$ for the binary case.
So far, we assumed that $\Phi(\Gamma)$ takes values in a countable set, that is, $P(\Phi)$ is a discrete distribution. Consider the case where $\Phi(\Gamma)$ produces continuous values, where the summation should be replaced by an integration. When considering a continuous distribution, the notable difference is that the differential entropy may take negative values. However, the existence of negative values is not problematic, as such negative values are offset in $H[P(\Phi)]-H[P(|\Phi|)]$. Therefore, the main result of Eq. (10) is well defined for the continuous case as well.
The derivation of Eq. (10) assumed that the observable $\Phi(\Gamma)$ does not include $\Phi(\Gamma)=0$. It is straightforward to extend the result to the case where the observable includes $\Phi(\Gamma)=0$. Specifically, when $\Phi(\Gamma)$ can equal 00, the range of $\Lambda[P(\Phi)]$ is modified as follows:
@@
which includes Eq. (10) as the specific case $P(\Phi=0)=0$. As long as the entropy production is given by Eq. (6), Eq. (14) holds for any observable $\Phi(\Gamma)$ satisfying Eq. (7). In Eq. (14), $\Phi=0$ plays a special role. When considering the process of doing nothing, the observable $\Phi$ is always 00 implying $P(\Phi=0)=1$. For such empty dynamics, the entropy production is 00, showing that both sides of Eq. (14) equal 00 and thus the inequality becomes an equality. However, note that this exceptional handling of $\Phi=0$ is not limited to Eq. (14). In the conventional thermodynamic uncertainty relation [Eq. (1)], when the expectation of the current vanishes, the inequality is ill-defined. Equations (10) and (14) saturate if and only if the system in equilibrium. For a detailed discussion on the saturation condition, see the supplementary material [26]. Again, when $\Phi(\Gamma)$ is a discrete distribution, the following relation holds as well:
$\Sigma\geq[1-P(\Phi=0)]\ln 2-H[P(\Phi)].$ (15)
@@
We remark on the relation between Eq. (14) and the Landauer principle [27]. Consider a process that resets a system, which has an equal probability of being in one of two states, to a single specific state. The following relation is obtained:
$\Delta S_{\mathrm{m}}\geq\ln 2,$ (16)